OpenStudy (anonymous):

Hi everyone! Can someone tell me why my teacher did this: |x-9|<1 so then -1

OpenStudy (anonymous):

is it because the limit of |x-9| as "n" goes to infinity is simply |x-9|? I just don't see why you just drop the absolute value signs once you put it into an inequality?

OpenStudy (anonymous):

what do you know when you have something in absolute?

OpenStudy (anonymous):

i mean..what is the definition of the absolute value?

OpenStudy (raden):

because for term in absolut is x-9, not x+9 :)

OpenStudy (raden):

the solution of ineq : |f(x)| < c, (with c positive) is -c < f(x) < c

OpenStudy (anonymous):

\[\left| x-9 \right|=\] x-9 if x>0 -(x-9) if x<0 for those 2 solutions if you put them into your equations you will get x<10 and x>8

OpenStudy (anonymous):

sorry x-9 if x-9>0 and -(x-9) if x-9<0

OpenStudy (anonymous):

not sure i follow you...any chance you can answer why he removed the absolute value sign but left it x-9 instead of changing it to x+9 using english rather than math examples? I just need the intuition here why it wasn't changed to x+9...thanks

OpenStudy (amistre64):

let k = x-9 |k| < 1, implies that: -1 < k < 1 since k = x-9; -1 < x-9 < 1 you appear to be confusing the value of (x-9) with the visual clue that an absolute value suggests. an absolute value function does not turn all subtraction signs into addition signs ....

OpenStudy (anonymous):

ok...i get it now! thanks everyone!

OpenStudy (amistre64):

:) good luck